Of the 23 homes in a subdivision, 7 need roofing and 5 need landscaping. If two homes need both roofing and landscaping, how many need neither?
Found 2 solutions by Hairydpm, jim_thompson5910:
Answer by Hairydpm(1)
Answer by jim_thompson5910(35256)

There are 7-2 = 5 homes that need only roofing

There are 5-2 = 3 homes that need only landscaping

So there are 5+3+2 = 10 homes that need at least one of the two

Therefore, there are 23 - 10 = 13 homes that need neither.